OpenStudy (anonymous):

Please Help!!!! I will give a medal and become a fan!!! 1. Use Euler's formula to find the missing number. Vertices:13 Edges:26 Faces:? (1 point)

OpenStudy (amistre64):

eulers formula is just basic addition ... whats the difficulty?

OpenStudy (anonymous):

I really need to get the right answer because i need to bring my grade up a lot

OpenStudy (amistre64):

well, then you need to apply eulers formula. and then we can see if you have the right answer. This site is not a free answering service.

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

hmm, first things first; can you show me you know eulers formula? if so then I can verify your results or correct it as needed

OpenStudy (anonymous):

\[e{ix} = \cos x + i \sin x\]

OpenStudy (amistre64):

thats not going to be it

OpenStudy (amistre64):

that is ONE of eulers things, but not quite the one we are looking for

OpenStudy (amistre64):

lets try: verts + faces = edges + 2

OpenStudy (amistre64):

like i said, simple addition :) 13 + f = 26 + 2 f = 26+2-13

OpenStudy (anonymous):

15

OpenStudy (amistre64):

15 is what i would go with yes.